# SciCoQA: Quality Assurance for Scientific Paper–Code Alignment

### v1

@@ 6 Conclusion
Our real-world data is predominantly skewed towards Computer Science and Artificial Intelligence. While we mitigate this by including synthetic discrepancies from Physics, Quantitative Biology, and Electrical Engineering, the distribution of errors in these fields may differ from our synthetic approximations. Consequently, model performance on non-CS domains should be interpreted with this synthetic nature in mind.
@@
With 611 discrepancies, our dataset is relatively small compared to large-scale pre-training corpora. This size is a deliberate trade-off to ensure high quality: real-world discrepancies are naturally sparse, and we employed a rigorous manual verification process to guarantee that every entry constitutes a meaningful mismatch rather than a trivial error or noise.
Ethical Considerations
@@
We constructed the synthetic portion of SciCoQA using repositories with permissive licenses (MIT, Apache 2.0, BSD, CC-BY) to ensure respectful redistribution of code. For real-world discrepancies derived from GitHub issues and reproducibility reports, we leverage the publicly available data to create our ground truth data by contextualizing and rephrasing the original issue. None of our data contains any personal information; however, since we work with scientific publications, these are closely associated with the authors of the respective papers. We acknowledge that highlighting discrepancies in a specific author’s work may be perceived negatively. We emphasize that these discrepancies are treated as scientific artifacts for improving community reproducibility standards, rather than criticisms of individual researchers.
@@
The benchmarked models in this work are intended to assist in quality assurance of scientific papers and their codebases. However, there is a risk of over-reliance; given the low recall rates demonstrated in our experiments ($45.7\%$ for GPT-5), automated tools should not yet be used as the sole arbiter of a paper’s validity. Relying blindly on these systems could lead to a false sense of security regarding the reproducibility and validity of a paper.
@@ Acknowledgments
Appendix A GitHub Crawl Details
@@
Appendix B Prompts
B.1 GitHub Issue Discrepancy Extraction
GitHub Issue Discrepancy Extraction You are an assistant that analyzes GitHub issues of scientific codebases. Your primary goal is to determine if the code repository contains any inconsistency, discrepancy, or mismatch between what is described in the paper and implemented in the code. For this, you analyze a GitHub issue and determine whether it reports a concrete discrepancy. A **concrete discrepancy** means the issue clearly describes a mismatch between what is stated in the paper (e.g., formulas, algorithms, hyperparameters, methods, logic, or processes) and what is implemented in the repository’s code. Important: Only label issues as a discrepancy if they point to a _specific, concrete difference_ between paper and code. Do not label general reproducibility problems, missing details, or unrelated bugs as discrepancies. ## Not a discrepancy issue Label as **Not a discrepancy issue** if: - The issue is about anything other than a difference between the paper and the code. - The issue describes reproducibility problems (e.g., different results) but does **not** identify a concrete paper-code mismatch. - The issue is about missing information needed to reproduce results without pointing to a mismatch. - The issue is about bugs or errors unrelated to the paper’s described methods or experiments. ## Discrepancy issue Label as **Discrepancy issue** if: - The issue explicitly reports a mismatch between the paper and the code implementation. - The mismatch can involve hyperparameters, formulas, algorithms, logic, processes, or other settings described in the paper. ## Response Format After analyzing the issue in detail and applying the definitions above, provide your final classification in the structure defined below: ### Final Answer “‘yaml issue_label: <the issue label: "Not a discrepancy issue" or "Discrepancy issue"> “‘ ## Issue {issue}
B.2 GitHub Discrepancy Verification
GitHub Issue Discrepancy Verification Your task is to verify a paper-code discrepancy described in a GitHub issue. Your goal is to verify whether the discrepancy is valid or not. Follow these steps to verify the discrepancy: 1. Analyze the issue and ensure you understand exactly the claimed discrepancy between the paper and the code. 2. Analyze the paper and the code and understand in detail the relevant paper sections and code files. 3. Using your understanding of paper, code, and the discrepancy in the issue, analyze whether the discrepancy is valid or not. A discrepancy is valid, if the claimed discrepancy actually exists and there is a difference between the paper description and the code. 4. Provide your final judgement whether the reported discrepancy is valid or not, and if so a summary and the relevant paper sections and code files in the format below: “‘yaml is_valid_discrepancy: <yes or no> is_valid_discrepancy_reason: <provide a short explanation for your judgement> discrepancy_summary: <if valid provide the following description, else this should be empty: a summary of the discrepancy between the paper and the code in 3-8 sentences. Your description should contain three parts focusing on the discrepancy: 1) summarize what is described in the paper, 2) summarize what is implemented in the code, and 3) summarize the difference. Do not speculate about the impact.> relevant_paper_sections: - <verbatim quote any parts from the paper that are relevant to the discrepancy> - <if there are multiple relevant parts, paste each of them.> relevant_code_files: - <name any code files that are relevant to the discrepancy by providing the file name> - <if there are multiple relevant code files, paste each of them.> “‘ ## Issue {issue} ## Paper {paper} ## Code {code}
B.3 Reproducibility Paper Discrepancy Extraction
Reproducibility Paper Discrepancy Extraction You are an assistant that analyzes reproducibility reports of scientific papers. Your primary goal is to detect whether the report identifies any discrepancies between the original paper and the original code repository. ## What counts as a discrepancy - A concrete paper–code discrepancy means the report clearly describes a mismatch between what is stated in the original paper (e.g., formulas, algorithms, logic, methods, processes, or other settings) and what is implemented in the original code repository. - Each distinct mismatch should be reported as a separate item. - Hyperparameter mismatches (e.g., learning rate, batch size, dropout rate) do not count as discrepancies, since these are typically configurable in code repositories. - If the report only describes differences between reproduced results and original results, without identifying a paper–code mismatch, do not list it. - If the report speculates about a possible mismatch (uncertain or ambiguous wording), still list it, but mark it with confindence: low. ## What does not count as a discrepancy - General reproducibility problems (e.g., “we could not match the reported results”). - Missing information in the paper (e.g., “the authors did not specify X”). - Missing implementation in the original code repository (e.g., “the authors did not provide the code for X”). - Bugs or errors in the code that are unrelated to what the paper describes. - Differences between reproduced implementation/results and the original paper. ## Output format Summarize all discrepancies found in the following structure, providing a description of the discrepancy, evidence supporting the discrepancy as verbatim quotes from the reproducibility report, and a confidence estimate of the authors on the reported discrepancy. “‘yaml concrete_paper_code_discrepancies: - description: "<3–8 sentence descriptive summary of the first discrepancy>" evidence: - "<paste any evidence (e.g. a paragraph describing the discrepancy) from the reproducibility report that support the discrepancy.>" - "<If there are multiple evidence, paste each of them.>" confidence: <Estimate of the confidence the authors have on the reported discrepancy. One of: low, medium, high>- … “‘ If no discrepancies are reported, return: “‘yaml concrete_paper_code_discrepancies: [] “‘ Reproducibility Report: {paper}
B.4 Reproducibility Paper Discrepancy Verification
Reproducibility Paper Discrepancy Verification Your task is to verify a paper-code discrepancy described in a reproducibility report. You will be provided with a summary of the discrepancy according to the report, and potentially multiple quotes from the reproducibility report that support the discrepancy. Your goal is to verify whether the discrepancy is valid or not. Follow these steps to verify the discrepancy: 1. Analyze the description and evidence from the reproducibility report and ensure you understand exactly the claimed discrepancy between the paper and the code. 2. Analyze the paper and the code and understand in detail the relevant paper sections and code files. 3. Using your understanding of paper, code, and the reported discrepancy, analyze whether the discrepancy is valid or not. A discrepancy is valid, if the claimed discrepancy actually exists, i.e. the described difference between paper and code exists. 4. Provide your final judgement whether the reported discrepancy is valid or not, and if so a summary and the relevant paper sections and code files in the format below: “‘yaml is_valid_discrepancy: <yes or no> is_valid_discrepancy_reason: <provide a short explanation for your judgement> discrepancy_summary: <if valid provide the following description, else this should be empty: a summary of the discrepancy between the paper and the code in 3-8 sentences. Your description should contain three parts focusing on the discrepancy: 1) summarize what is described in the paper, 2) summarize what is implemented in the code, and 3) summarize the difference. Do not speculate about the impact.> relevant_paper_sections: - <verbatim quote any parts from the paper that are relevant to the discrepancy> - <if there are multiple relevant parts, paste each of them.> relevant_code_files: - <name any code files that are relevant to the discrepancy by providing the file name> - <if there are multiple relevant code files, paste each of them.> “‘ ## Reproducibility Report with Discrepancy {discrepancy_in_report} ## Paper {paper} ## Code {code}
B.5 Synthetic Discrepancy Generation
Synthetic Discrepancy Generation Your task is to generate 5 realistic paper–code discrepancies by introducing small, conceptually meaningful modifications to the codebase of the provided research paper. Follow these steps to perform the edit: 1. Carefully read and understand both the research paper and the entire code repository provided below. Your goal is to identify the key ideas, methods, and components described in the paper and how they correspond to the implementation in the code. 2. Your changes must adhere to the following constraints: - **Small**: The changes must affect a few lines of code or a short function. It may affect multiple files, but only to the extent necessary to create a coherent and realistic discrepancy. - **Relevance**: The changes must relate directly to a core scientific or algorithmic idea of the paper and would likely have an impact on reproducibility of results and validity of claims. - **Significance**: The changes must introduce a conceptual discrepancy — not a simple hyperparameter, or formatting change, or a change that could be fixed via a simple (command line) argument. - **Scope**: The changes should not all rely on already implemented features of the code base, but also implement new features or introduce modifications. Balance between relying on existing features and implementing new features. - **No Comments**: Do not add comments to the changed code which would easily allow to identify the discrepancy. - **No Bugs**: The introduced discrepancies should not be bugs, i.e., code that could be detected as erroneous by inspecting the code itself. The discrepancy must be related to both the paper and the code to be identified. - **No Implications**: The discrepancies must not rely on anything the paper only implies or assumes, but the paper must clearly conflict with the code after the change, or omit an important concept that is implemented in the changed code. 3. Your changes can be one of the following types. You can create multiple discrepancies of the same or different types. - **Paper Omission**: Modify the code such that it implements a concept or idea that is not described in the paper. - **Code Omission**: Modify the code such that it drops a specific concept or idea that is described in the paper. - **Difference**: Modify the code such that there is a difference between the paper and the code, i.e., the paper describes one thing and the code implements another, e.g., by changing the order of steps, operations, or core logic. 4. Decide which paper-code discrepancies are most appropriate for the given paper, choosing from the following categories. Note, you can create multiple discrepancies of the same or different types. - **Loss**: changes to loss definition or terms - **Algorithm**: changes in the order of steps, operations, or core logic - **Training**: changes to the learning process, schedule, or optimization - **Evaluation**: changes to evaluation logic, metrics, or scripts - **Model**: architectural or initialization changes - **Data**: dataset usage, preprocessing, augmentation, or filtering - **Other**: other types of discrepancies that are not covered by the above categories but appropriate for the given paper. 5. Generate 5 discrepancies in the following strict format: “‘md # Discrepancy 1 - Type: <choose one from: Paper Omission, Code Omission, Difference> - Category: <choose one from: Loss, Algorithm, Training, Evaluation, Model, Data, Other> - Description: <a summary of the discrepancy between the paper and the code in 3-8 sentences. When referring to the code, do not mention ’original’ or ’modified’, but assume the code is published with the modification and discrepancy. Your description should contain three parts focusing on the discrepancy: 1) summarize what is described in the paper, 2) summarize what is implemented in the modified code, and 3) summarize the difference. Do not speculate about the impact.> ## Code Changes <<<<<<<ORIGINAL CODE: <relative/path/to/file.py> (paste the relevant lines of the original code exactly as they appear) ======= (paste the lines of the modified code containing your change that replace the original code) >>>>>>>DISCREPANCY If multiple files are affected, repeat the diff block for each. ## Relevant Paper Sections - <verbatim quote any parts from the paper that are relevant to the discrepancy.> - <if there are multiple relevant parts, paste each of them.> # Discrepancy 2 # Discrepancy 3 “‘ ## Paper {paper} ## Code {code}
B.6 Discrepancy Prediction
Discrepancy Prediction You are an expert in analyzing scientific papers and their code implementations. Your task is to carefully identify concrete discrepancies between what is described in a paper and what is actually implemented in the code. ## What counts as a discrepancy - A concrete paper–code discrepancy means a mismatch between what is stated in the original paper (e.g., formulas, algorithms, logic, methods, processes, or other settings) and what is implemented in the original code repository. - Each distinct mismatch should be reported as a separate item. ## What does not count as a discrepancy - Missing information in the paper like hyperparameters (e.g., “the authors did not specify X”). - Hyperparameter mismatches (e.g., learning rate, batch size, dropout rate), since these are typically configurable in code repository. - Missing implementation in the original code repository (e.g., “the authors did not provide the code for X”). - Bugs or errors in the code that are unrelated to what the paper describes. ## Output format Provide your findings in the following YAML structure: “‘yaml discrepancies: - <a summary of the discrepancy between the paper and the code in 3-8 sentences. Your description should contain three parts focusing on the discrepancy: 1) summarize what is described in the paper, 2) summarize what is implemented in the code, and 3) summarize the difference. Do not speculate about the impact.> - <if there are multiple discrepancies, put each of them in a separate item.> “‘ ## Paper {paper} ## Code
B.7 Discrepancy Evaluation
Discrepancy Evaluation Your task is to evaluate whether a reference paper - code discrepancies matches a predicted paper - code discrepancy. Follow these steps: 1. Analyze which part of the paper or code each discrepancy is describing. Extract the core claims and issues from the reference and predicted discrepancies. 2. Analyze whether the core claims are about the same issue, i.e. if they describe the same or different paper-code discrepancies. The two discrepancies might use different wording or one might be more detailed than the other. Focus on whether the issue is the same, even if minor details are different. However, if they describe different issues (even about the same topic or part of the paper or code) they do not match. 3. Provide a brief explanation of your reasoning. ## Reference Paper-Code Discrepancy {reference discrepancy} ## Predicted Paper-Code Discrepancy {predicted discrepancy} ## Answer Format Provide your answer in the following format: “‘yaml core_claim_reference: <core claim from reference discrepancy > core_claim_predicted: <core claim from predicted discrepancy > reasoning: <explanation of why the core claims concern the same issue > match: <yes | no > “‘
Appendix C Implementation Details
C.1 Issue Processing
@@ C.2 Versioning
C.3 Paper Processing
We provide as input the paper text in markdown. We use MistralOCR [^6] https://mistral.ai/news/mistral-ocr to convert the PDF to markdown. We exclude figures, although the captions are included. We further remove the references section from all papers.
@@ C.4 Code Processing
We filter the files to include only files that contain relevant content, mainly excluding dataset files. Specifically, we include files with the following extensions: .c, .cc, .cpp, .cu, .h, .hpp, .java, .jl, .m, .matlab, .Makefile, .md, .pl, .ps1, .py, .r, .sh, config.txt, .rs, readme.txt, requirements_dev.txt, requirements-dev.txt, requirements.dev.txt, requirements.txt, .scala, .yaml, .yml. We obtain this list by manually inspecting the file extensions in the repositories of our dataset. We further process all jupyter notebook files to Python files by removing their output. For the prompt, we further include a list of all files in the repository to the model.
C.5 Context Processing
@@
C.6 Model and Decoding Configuration
@@
Appendix D Synthetic Code Analysis
We conduct a quantitative analysis of the synthetic data to validate that our generation pipeline produces minimal changes, leaving the main code intact. Figure 6 presents the analysis of the generated code diffs.
@@ Appendix D Synthetic Code Analysis
We find the generated discrepancies in the code to be highly targeted. The vast majority affect only a single file ($1.08\pm 0.29$ on average), with a maximum of three files. Similarly, the number of distinct generated diffs is low, with a mean of $1.24\pm 0.59$ snippets per discrepancy. This confirms that the generated errors are specific deviations in logic rather than large-scale refactors. Similarly, the number of line edits (i.e., added or removed lines) further supports the subtlety of the dataset. The distribution of line counts is heavily skewed towards small edits: On average, $2.31\pm 2.78$ lines are added and $3.17\pm 4.51$ lines are removed. The slightly higher count for removed lines suggests the discrepancies often simplify logic (e.g., removing a normalization step) rather than adding complex boilerplate.
@@
To quantify the similarity between the original and modified code, we calculate the Jaccard similarity. We measure the similarity between the entire generated code (which often contains a few more lines that are not changed), and only code lines that were modified. Comparing the entire code block generated by the model against the original yields a high mean similarity of $0.74\pm 0.20$. This indicates that the surrounding context and structure remain largely identical. Comparing only the specific lines that differ (added vs. removed) yields a mean similarity of $0.50\pm 0.29$. This moderate overlap suggests the changes are often modifications of existing statements (e.g., changing an operator or variable) rather than complete replacements. Overall, the data confirms that the synthetic discrepancies are precise and minimal, strictly adhering to the “small” and “conceptually meaningful” constraints of the synthetic data generation prompt (§B.5).
Appendix E Analysis by Programming Language
Figure 7 shows the distribution of programming languages in the SciCoQA dataset. Since the real data originates from the CS/AI domain, we observe Python as the dominant language, along with a small portion of C/C++, Matlab, and CUDA files. Besides Python, other languages are present in the synthetic data, including Java, Scala, Julia, and R, although their overall presence is relatively small.
@@ Appendix E Analysis by Programming Language
E.1 Results by Programming Language
In the synthetic dataset, the known injection points allow us to analyze detection performance across different programming languages. Since the dataset is skewed towards Python, drawing conclusions for specific low-resource languages is difficult due to the small sample size. Therefore, we group all non-Python instances into a unified subset of $46$ samples for analysis ("Not Python"). The resulting recall scores are presented in Table 4

### v2

@@ 7 Limitations
Our real-world data is predominantly skewed towards Computer Science and Artificial Intelligence. While we mitigate this by including synthetic discrepancies from Physics, Quantitative Biology, and Electrical Engineering, the distribution of errors in these fields may differ from our synthetic approximations. Consequently, model performance on non-CS domains should be interpreted with this synthetic nature in mind. We applied our real-world pipeline also to non-CS domains, searching GitHub repositories for URLs linking to PubMed, bioRxiv, medRxiv, but found very few with associated issues reporting discrepancies. Further, to our knowledge, no domain outside CS/AI systematically produces reproducibility papers with paper-code discrepancies. For this reason, we expanded to other domains via synthetic data, which we believe provides more value than excluding these domains entirely and we include a call for more reproducibility efforts in the conclusion.
@@
With $635$ discrepancies, our dataset is relatively small compared to large-scale pre-training corpora. This size is a deliberate trade-off to ensure high quality: real-world discrepancies are naturally sparsely documented, and we employed a rigorous manual verification process to guarantee that every entry constitutes a meaningful mismatch rather than a trivial error or noise.
Synthetic GPT-5 Performance
Our synthetic discrepancies are generated by GPT-5, which is also evaluated on the benchmark. We find that GPT-5 achieves disproportionately higher recall on synthetic data. This suggests that GPT-5 may benefit from its involvement in the data generation on this subset. However, the strong correlation between real and synthetic recall across all $22$ models ($r=0.94$, Fig. 9) confirms that relative model comparisons remain valid. Notably, excluding the GPT-5 family from the correlation analysis increases the correlation to $r=0.98$, indicating that GPT-5’s synthetic advantage is an outlier that slightly weakens the overall trend rather than inflating it. To mitigate this limitation, absolute model performance should be assessed on the real subset, while the synthetic subset reliably supports relative model comparison and allows for targeted generation of new, uncontaminated data.
8 Ethical Considerations
@@
We constructed the synthetic portion of SciCoQA using repositories with permissive licenses (MIT, Apache 2.0, BSD, CC-BY) to ensure respectful redistribution of code. For real-world discrepancies derived from GitHub issues and reproducibility reports, we leverage the publicly available data to create our ground truth data by contextualizing and rephrasing the original issue. None of our data contains any personal information; however, since we work with scientific publications, these are closely associated with the authors of the respective papers. We acknowledge that highlighting discrepancies in a specific author’s work may be perceived negatively. We emphasize that these discrepancies are treated as scientific artifacts for improving community reproducibility standards, rather than criticisms of individual researchers. Discrepancies frequently arise from benign causes, including concurrent code updates, simplifications for readability, or distinct authorship of paper and code, and their presence does not imply negligence or misconduct.
@@
The benchmarked models in this work are intended to assist in quality assurance of scientific papers and their codebases. However, there is a risk of over-reliance; given the low recall rates demonstrated in our experiments ($46.7\%$ for Gemini 3.1 Pro and GPT-5 Mini), automated tools should not yet be used as the sole arbiter of a paper’s validity. Relying blindly on these systems could lead to a false sense of security regarding the reproducibility and validity of a paper. We emphasize that model outputs should support human review and reproducibility efforts, not serve as automatic reject signals or as evidence of bad faith. Automated detections, particularly given the false positive rates observed in our experiments, require expert verification before any conclusions about research quality are drawn.
@@ Acknowledgments
Appendix A Implementation Details
A.1 GitHub Crawl Details
@@
A.2 Issue Processing
@@ A.3 Versioning
A.4 Paper Processing
We provide as input the paper text in markdown. We use Mistrals OCR (Mistral AI Team, 2025) to convert the PDF to markdown. We exclude figures, although the captions are included. We further remove the references section from all papers.
@@ A.5 Code Processing
We filter the files to include only files that contain relevant content, mainly excluding dataset files. Specifically, we exclude files that are larger than 1MB and include files with the following extensions: .c, .cc, .cpp, .cu, .h, .hpp, .java, .jl, .m, .matlab, Makefile, .md, .pl, .ps1, .py, .r, .sh, config.txt, .rs, readme.txt, requirements_dev.txt, requirements-dev.txt, requirements.dev.txt, requirements.txt, .scala, .yaml, .yml. We obtain this list by manually inspecting the file extensions in the repositories of our dataset. We further process all jupyter notebook files to Python files by removing their output. For the prompt, we further include a list of all files in the repository to the model.
A.6 Context Processing
@@
A.7 Model and Decoding Configuration
@@
Appendix B Extended Analysis
B.1 Gemini and GPT Ground Truth Comparison
Figure 7: Comparisons between the two ground truth descriptions generated by Gemini 3.1 Pro and GPT-5.
For the real data we rephrased the ground truth discrepancy descriptions with Gemini 3.1 Pro and GPT-5 (see §3.1). We analyze the similarity between these two descriptions in Fig. 7. We find the GPT descriptions to be slightly longer ($181\pm 40$ vs $135\pm 34$ tokens). We further find moderate lexical overlap between the two ground truths (average TF-IDF cosine similarity is $0.52$), but high semantic similarity (average $0.83$ computed using GemmaEmbedding), indicating that the models use different wording to describe the same discrepancy. Finally, we analyze how often the GPT-OSS 20B judge matches a prediction against none, either or both ground truths. We find that in $97.7\%$, the judge comes to the same verdict, given the two different discrepancy descriptions, yielding a Cohen’s Kappa of $\kappa=0.83$.
B.2 Synthetic Code Analysis
We conduct a quantitative analysis of the synthetic data to validate that our generation pipeline produces minimal changes, leaving the main code intact. Fig. 6 presents the analysis of the generated code diffs.
@@ B.2 Synthetic Code Analysis
We find the generated discrepancies in the code to be highly targeted. The vast majority affect only a single file ($1.08\pm 0.29$ on average), with a maximum of three files. Similarly, the number of distinct generated diffs is low, with a mean of $1.24\pm 0.61$ snippets per discrepancy. This confirms that the generated errors are specific deviations in logic rather than large-scale refactors. Similarly, the number of line edits (i.e., added or removed lines) further supports the subtlety of the dataset. The distribution of line counts is heavily skewed towards small edits: On average, $2.41\pm 3.25$ lines are added and $3.13\pm 4.47$ lines are removed. The slightly higher count for removed lines suggests the discrepancies often simplify logic (e.g., removing a normalization step) rather than adding complex boilerplate.
@@
To quantify the similarity between the original and modified code, we calculate the Jaccard similarity. We measure the similarity between the entire generated code (which often contains a few more lines that are not changed), and only code lines that were modified. Comparing the entire code block generated by the model against the original yields a high mean similarity of $0.74\pm 0.20$. This indicates that the surrounding context and structure remain largely identical. Comparing only the specific lines that differ (added vs. removed) yields a mean similarity of $0.50\pm 0.29$. This moderate overlap suggests the changes are often modifications of existing statements (e.g., changing an operator or variable) rather than complete replacements. Overall, the data confirms that the synthetic discrepancies are precise and minimal, strictly adhering to the “small” and “conceptually meaningful” constraints of the synthetic data generation prompt (§F.5).
B.3 Programming Language Distribution
Fig. 8 shows the distribution of programming languages in the SciCoQA dataset. Since the real data originates from the CS/AI domain, we observe Python as the dominant language, along with a small portion of C/C++, Matlab, and CUDA files. Besides Python, other languages are present in the synthetic data, including Java, Scala, Julia, and R, although their overall presence is relatively small.
@@ B.3 Programming Language Distribution
Appendix C Extended Results
C.1 Error Analysis
Model GPT-5 Gemini GPT-OSS
False Positives $10$ 33 $34$
Code Misunderstanding 11 0 2020
Paper Misunderstanding 22 0 99
Paper Ambiguity 22 11 0
OCR Error 11 11 11
3rd Party Code 22 0 11
No Discrepancy 11 0 11
Minor (Config) 11 0 22
Other 0 11 0
True Positives $60$ $48$ $69$
Unique $23$ 1919 $34$
Minor 77 77 $12$
Total $70$ $51$ $103$
Precision $85.7$ $94.1$ $67.0$
Table 3: Analysis of unlabeled discrepancy predictions of GPT-5, Gemini 2.5 Pro, and GPT-OSS 20B on 2020 NLP and CV papers from the real subset of SciCoQA.
Table 3 provides the detailed breakdown of our human analysis from §5.2 on the predictions of GPT-5, Gemini 2.5 Pro and GPT-OSS 20B.
C.2 Performance by Programming Language
In the synthetic dataset, the known injection points allow us to analyze detection performance across different programming languages. Since the dataset is skewed towards Python, drawing conclusions for specific low-resource languages is difficult due to the small sample size. Therefore, we group all non-Python instances into a unified subset of $46$ samples for analysis (“Not Python”). The resulting recall scores are presented in Table 4